if twice the son's age, in years, is added to the father's age, the sum is 70. but if twice the father's age is added to the son's age the sum is 95. find the ages of the father and the son.
Answer by stanbon(75887)

if twice the son's age, in years, is added to the father's age, the sum is 70. but if twice the father's age is added to the son's age the sum is 95. find the ages of the father and the son.
-----------------------------------------------
Equations:
2s + f = 70
s + 2f = 95
------------------------
Multiply thru the 1st equation by 2 and solve for "s":
4s + 2f = 140
s + 2f = 95
----------------------
Subtract 2nd from 1st:
3s = 45
s = 15 yrs (son's age)
------------------
Substitute into 2s+f = 70 to solve for "f":
2*15 + f = 70
f = 40 (father's age)
===========================
Cheers,
Stan H.